during a rainy day ,rain is falling vertically with a velocity 2m/s. a boy at rest starts his motion with a constant acceleration of 2m/s2 along a straight road. is the rate at which the axis of umbrella with vertical should be changed is 1/n at t=5 sec so that the rain always falls parallel to the axis of the umbrella , then find n.
 

Grade:12th pass

2 Answers

Sandeep
6 years ago
horizontally component of V(rain velocity):2sinAfor man:u=0; v=0+at 2sinA=2t......... (1) cosA(dA/dt)=1 dA/dt=secANow, time is given so we can get the value of A from (1) and put in above eqn and get da/dt
Rituljangir
5 years ago
Velocity of man after time t is 
V= 2t 
Velocity of rain is 2 
Tan a = velocity of man / rain 
Tan a = t
Diff this eq with respect to t
Sec^ 2 a da/ dt =1 
(Tan^ 2 a +1) da/ dt = 1 
da / dt = 1/ t^ 2 + 1